Your hotel tells you that they will increase the temperature by 10 percent each hour. the current temperature of the hot tub is 75 degrees F.
Write out the first 3 terms of the sequence, rounding to the nearest hundredth

Answer by stanbon(75887)

Your hotel tells you that they will increase the temperature by 10 percent each hour. the current temperature of the hot tub is 75 degrees F.
Write out the first 3 terms of the sequence, rounding to the nearest hundredth
--------------------
75
75*(1.10) = 82.5
75*(1.10)^2 = 90.75
75*(1.10)^3 = 99.825
===============
Cheers,
Stan H.
---------------